OpenStudy (anonymous):

Find the line which is perpendicular to the graph of fx=x/(x+2) at x=2

OpenStudy (anonymous):

find derivative, find slope

OpenStudy (anonymous):

you can first find y by plugging in x

OpenStudy (anonymous):

y=2/4

OpenStudy (anonymous):

now you have x,y need slope using derivative and then you are set to go

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

f'(x)=((x+2)(1)+(x)(1))/(x+2)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

umm i use product rule give me a sec lol

OpenStudy (anonymous):

f'(x) = 2/(2+x)^2 slope: 2/(2+x)^2 slope 2/(2+2)^2 slope = 1/8 perpendicular slope: -8 f(2) = 2/4 = 1/2 y = -8x +b 1/2 = -8 (2) +b b = 9/2 y = -8x + 9/2

OpenStudy (anonymous):

oops i mean - not +

OpenStudy (anonymous):

i got -8/3 as my slope and my y=-8/3x+(35/6) is that even close?